OpenStudy (anonymous):

4x^2+33x+35 what is the factored form ?

OpenStudy (saifoo.khan):

srsly?

OpenStudy (anonymous):

- Do I juss factor the middle & last term ? :/

OpenStudy (saifoo.khan):

(x+7) (4 x+5)

OpenStudy (anonymous):

4x^(2)+33x+35 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (140) that add up to b (33).In this problem 7*(5)/(4)=(35)/(4) (which is (c)/(a)) and 7+(5)/(4)=(33)/(4) (which is ((b)/(a)) , so insert 7 as the right hand term of one factor and (5)/(4) as the right-hand term of the other factor. (x+7)(x+(5)/(4)) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (x+7)(4x+5)